# Generative Diffusion Model Driven Massive Random Access in Massive MIMO Systems

### v2

@@ Abstract
Massive random access is an important technology for achieving ultra-massive connectivity in next-generation wireless communication systems. It aims to address key challenges during the initial access phase, including active user detection (AUD), channel estimation (CE), and data detection (DD). This paper examines massive access in massive multiple-input multiple-output (MIMO) systems, where deep learning is used to tackle the challenging AUD, CE, and DD functions. First, we introduce a Transformer-AUD scheme tailored for variable pilot-length access. This approach integrates pilot length information and a spatial correlation module into a Transformer-based detector, enabling a single model to generalize across various pilot lengths and antenna numbers. Next, we propose a generative diffusion model (GDM)-driven iterative CE and DD framework. The GDM employs a score function to capture the posterior distributions of massive MIMO channels and data symbols. Part of the score function is learned from the channel dataset via neural networks, while the remaining score component is derived in a closed form by applying the symbol prior constellation distribution and known transmission model. Utilizing these posterior scores, we design an asynchronous alternating CE and DD framework that employs a predictor-corrector sampling technique to iteratively generate channel estimation and data detection results during the reverse diffusion process. Simulation results demonstrate that our proposed approaches significantly outperform baseline methods with respect to AUD, CE, and DD.
@@ I Introduction
In wireless communications, multi-user random access is a critical step in establishing initial communication links. This process involves techniques such as activity user detection (AUD), channel estimation (CE), and multi-user data detection (DD). Traditional cellular networks often use scheduling-based access protocols and orthogonal multiple access methods [1]. As networks evolve, the 6G vision outlines more stringent performance requirements for scenarios involving massive communication. This vision necessitates a connection density ranging from $10^{6}$ to $10^{8}$ devices per square kilometer[2]. As the number of users increases, dependence on orthogonal multiple access will lead to excessive overhead, substantially affecting communication efficiency.
To overcome the limitations of existing scheduling-based random access protocols, the grant-free non-orthogonal multiple access scheme has been proposed [1]. By eliminating the need for complex handshaking between users and the base station (BS), terminals can directly transmit uplink signals composed of non-orthogonal preamble sequences and payload data. This approach effectively reduces communication delays while enabling more devices to access the network simultaneously. However, it also increases the complexity of designing receiver detection algorithms. Previous studies have proposed several advanced receiver schemes, notably sourced random access and unsourced random access [3]. Sourced random access [4] employs pre-assigned preamble sequences to distinguish user identities and facilitates user activity detection and channel estimation using these preambles. In contrast, unsourced random access employs a common codebook, utilizing codewords to convey information and select coding, interleaving, and spreading techniques [5]. Moreover, reference [6] investigates a massive random access scheme utilizing individual user codebooks. This scheme recognizes the uniqueness of user codewords and encodes bits during the codeword selection process, integrating features from both sourced and unsourced random access mechanisms previously described. The reference [6] provides an in-depth performance analysis of massive random access under multiple-input multiple-output (MIMO) Rayleigh channels, taking into account the presence or absence of channel state information (CSI) at the receiver and whether the number of active users is known, thereby offering theoretical insights into performance limits. This paper mainly focuses on sourced random access schemes as in [4], tackling algorithm design challenges related to AUD, CE, and DD. For simplicity, we continue to refer to the discussed approach as massive random access.
@@ I-A Related Work
Once the AUD and CE processes are completed, the massive MIMO DD problem at BS often becomes overdetermined. Various advanced detection algorithms have been developed to address this challenge. In addition to traditional methods like least squares (LS) and linear minimum mean square error (LMMSE), algorithms such as OAMP [18], expectation propagation (EP) [19], and deep learning-based techniques [20, 21, 22, 23] have demonstrated excellent performance. However, these DD methods rely on the CSI estimated in the earlier stage, and errors in CSI estimation will degrade the detection performance.
Performing joint CE and DD (JCEDD) can enhance overall achievable performance. The work [21] presents an iterative JCEDD scheme that accounts for channel estimation errors during data detection and employs a model-driven OAMP unfolding algorithm to improve detection. In the CE phase, demodulated data are used as pilots to augment LMMSE CE, resulting in superior performance. Additionally, several studies have addressed the joint AUD, CE, and DD problems, where message-passing algorithms or optimization-based methods have been developed to tackle these tasks [24, 25, 26]. Despite their strong performance, these algorithms require elaborate system configurations and dedicated algorithm design, which may not generalize effectively to practical systems or complex channel environments that are challenging to model, thus limiting their applicability. Consequently, they are not included in the comparisons of this paper.
@@ I-B Our Contributions
This paper proposes a new two-stage algorithm to solve the AUD, CE, and DD tasks in massive random access systems. The first stage identifies the active users set (AUS) through a Transformer-based detector, while the second stage utilizes generative diffusion models (GDMs) to perform JCEDD for active user equipment (UEs). The existing Transformer-based deep learning AUD scheme [10] fails to effectively address the challenges posed by variable pilot lengths in massive random access problems. Furthermore, the performance of current JCEDD schemes relying on ALD-based generative methods [33] has not been comprehensively evaluated and is susceptible to converging to local optima. To tackle these issues, our research makes the following key contributions:
@@
  • We model the CE and DD tasks as reverse diffusion processes that sample from a joint distribution. These processes gradually transform initial noise into the original channel and data symbols. By employing advanced predictor-corrector samplers, our scheme outperforms existing JCEDD schemes.
  • To accommodate the distinct characteristics of unknown continuous channel distributions and known discrete data constellation distributions, we propose an asynchronous iterative CE and DD framework. This framework integrates an LMMSE-based CE initialization method, which reduces the number of sampling steps and enhances convergence performance.
@@ III Problem Formulation
The proposed massive random access system is depicted in Fig. 1, where the BS needs to solve the problems of AUD and JCEDD. Specifically, we divide the AUD, CE, and DD tasks into the two stages. In the first stage, we use a single model to derive the AUS information from the received pilots of variable lengths. In the second stage, we perform the JCEDD using the powerful generation capability of diffusion models.
@@ III-B Stage 2-JCEDD
Different from typical passing algorithms that factorize the joint distribution (8) into a factor graph and then inferring the posterior distribution of each element through message passing, our proposed scheme directly samples the signals from the posterior distributions (9a) and (9b). This approach was first implemented in [27, 31, 32, 33], where the ALD sampler is used to sample from the posterior distribution in the CE or DD problem.
@@ IV-B Heterogeneous Transformer and Activity Decoder
Assuming that the heterogeneous Transformer network consists of $N$ encoder layers, we conduct activity estimation for $K$ users based on the K+1K\!+\!1 sequences output by the $N$-th encoder layer. This task is performed by the activity decoder, which computes the heterogeneous MHA between the final output feature sequence of the received signal $\bm{z}_{K+1}^{N+1}$ and pilot sequences $\{\bm{z}_{k}^{N+1}\}_{k=1}^{K}$. The resulting MHA output, $\bm{z}_{c}$, represents the synthesized contributions of all pilot sequences to the received signal. Next, the user activity is estimated as the weighted correlation between $\bm{z}_{c}$ and each pilot feature sequence. It is given by:
@@ V Generative Diffusion Models Driven JCEDD Using Predictor-Corrector Sampler
Given the active UE set, we need to recover the channel and corresponding data symbols of the active UEs. Before introducing the proposed JCEDD schemes, we briefly review the relationship between existing generative models and their SDE descriptions. We then address the CE problem, detailing the calculation of posterior score functions within the predictor-corrector (PC) sampler framework. Subsequently, we extend the application of the PC sampler to the DD problem. Finally, we combine the CE and DD methods to propose an iterative algorithm for JCEDD and analyze its complexity.
@@ V-B3 Channel Sampling Process
Utilizing the derived prior score function and likelihood score function, we numerically solve the reverse SDE (V-B). Specifically, by dividing the interval $t\!\in\!(0,\,1]$ into $N$ discrete steps, indexed as $i\!=\!0,1,\dots,N-1$, we obtain the following iterative rule for the reverse diffusion sampler (which is also called predictor):
@@
where $\epsilon_{i}\!=\!2\sigma_{i}\big(r\|\bm{Z}\|_{2}/\|\nabla_{\tilde{\bm{H}}_{a}^{i}}\log p(\tilde{\bm{H}}_{a}^{i}|\bm{Y})\|_{2}\big)^{2}$ is the step size in the $i$-th iteration of the ALD algorithm [38]. Additionally, $r$ is an empirically chosen hyperparameter, set to $r=0.3$ in our simulations. Existing ALD-based studies on CE [27], DD [31], and JCEDD [33] can be regarded as corrector-only samplers with potential for performance improvement. Therefore, we address the JCEDD problem using the PC sampler illustrated in Fig. 4.
@@ V-C Data Detection
Upon reviewing the formulas (2) and (7), the DD problem can be characterized as $\bm{Y}_{d}\!=\!\bm{X}_{a}\bm{H}_{a}\!+\!\bm{W}_{d}$. By transposing the formula and representing the model in a real-valued form, the signal transmission model can be reformulated as:
@@
where $\lambda_{x}\!=\!2.5$ is an empirically chosen hyperparameter. The DD algorithm can also be derived based on PC-sampler. Due to space constraint, the detailed DD algorithm is omitted, but its basic iterative procedure is similar to that of Algorithm 1.
@@ V-D Overall Algorithm
In the CE and DD algorithms above, the data $\bm{X}_{a}$ is assumed to be known during CE, while the channel $\bm{H}_{a}$ is assumed to be known during DD. In practical transmission, only the pilot $\bm{P}_{a}$ and the received signals $\bm{Y}$ are known. Therefore, we design an iterative algorithm to perform JCEDD. The proposed PC sampler-based JCEDD method is depicted in Fig. 5, and the algorithmic framework is summarized as Algorithm 2.
The issues we need to address are: 1) A typical SDE sampling algorithm starts by sampling from random noise, resulting in a high number of iterations. 2) In the existing ALD-based JCEDD [33], each CE iteration is followed by one DD iteration. Due to the different distributions of the channel and data symbols, their convergence speeds are inconsistent, causing the DD to fall into local optima before the completion of the CE iterations. We address these issues by the following improvements in the proposed sampling mechanism.
@@
    By reshaping the estimated $\hat{\bm{h}}_{a}$ into matrix $\hat{\bm{H}}_{a}$, the initial DD result can be obtained by applying zero forcing (ZF) to (27), i.e., $\hat{\bm{X}}_{a}\!=\!\bm{Y}_{d}\bm{H}_{a}^{\rm H}\big(\bm{H}_{a}\bm{H}_{a}^{\rm H}\big)^{-1}$.
  • We design an asynchronous alternating iteration mechanism. For channel datasets with complex distributions, the calculation of posterior score is more complicated than DD that relies solely on model-based iterations with closed-form expression. We update the DD results after every $N_{\text{update}}$ channel iterations (Line 6 in Algorithm 2). To ensure the optimal DD results under the current channel error, we conduct $N_{X}$ steps of PC sampler each time the data symbols are updated (Lines 9-12).
@@
1:Initialize: Acquire initial CE $\hat{\bm{H}}_{a}$ via LMMSE, initial DD $\hat{\bm{X}}_{a}$ via ZF, and start point $i^{*}$ via (33);
@@ VI-A2 Channel Prior Score Network Setup
During the testing phase, the reverse SDE process for CE is discretized into $N_{H}\!=\!1,500$ steps. Given that an LMMSE initialization is employed, only $i^{*}$ of the $N_{H}$ steps are executed. The reverse SDE process for DD is also discretized into $N_{X}\!=\!1,500$ steps, using a noise scheduling function defined as $\tau(t)\!=\!\tau_{\min}\big(\frac{\tau_{\max}}{\tau_{\min}}\big)^{t}$, where $\tau_{\text{max}}\!=\!1$ and $\tau_{\text{min}}\!=\!0.01$. Additionally, the hyperparameters are set to $\lambda_{h}\!=\!\lambda_{x}\!=\!2.5$.
@@ VI-A3 Evaluation Metrics
We utilize activity error probability (AEP), normalized mean square error (NMSE) and bit error rate (BER) to evaluate the performance of AUD, CE and DD, respectively. These metrics are defined respectively as $\text{AEP}\!=\!\mathbb{E}\big\{\frac{1}{K}\sum_{k=1}^{K}|\alpha_{k}\!-\!\hat{\alpha}_{k}|\big\}$, $\text{NMSE}\!=\!\mathbb{E}\big\{\frac{\|\bm{H}_{a}\!-\!\hat{\bm{H}}_{a}\|^{2}}{\|\bm{H}_{a}\|^{2}}\big\}$ and $\text{BER}\!=\!\mathbb{E}\big\{\frac{E_{b}}{K_{a}L_{d}\log_{2}|\mathcal{X}|_{c}}\big\}$, where $E_{b}$ denotes the total number of error bits for active UEs.

### v3

@@ Abstract
Massive random access is an important technology for achieving ultra-massive connectivity in next-generation wireless communication systems. It aims to address key challenges during the initial access phase, including active user detection (AUD), channel estimation (CE), and data detection ($D$). This paper examines massive access in massive multiple-input multiple-output (MIMO) systems, where deep learning is used to tackle the challenging AUD, CE, and $D$ functions. First, we introduce a Transformer-AUD scheme tailored for variable pilot-length access. This approach integrates pilot length information and a spatial correlation module into a Transformer-based detector, enabling a single model to generalize across various pilot lengths and antenna numbers. Next, we propose a generative diffusion model (GDM)-driven iterative CE and $D$ framework. The GDM employs a score function to capture the posterior distributions of massive MIMO channels and data symbols. Part of the score function is learned from the channel dataset via neural networks, while the remaining score component is derived in a closed form by applying the symbol prior constellation distribution and known transmission model. Utilizing these posterior scores, we design an asynchronous alternating CE and $D$ framework that employs a predictor-corrector sampling technique to iteratively generate channel estimation and data detection results during the reverse diffusion process. Simulation results demonstrate that our proposed approaches significantly outperform baseline methods with respect to AUD, CE, and $D$.
@@ I Introduction
In wireless communications, multi-user random access is a critical step in establishing initial communication links. This process involves techniques such as activity user detection (AUD), channel estimation (CE), and multi-user data detection ($D$). Traditional cellular networks often use scheduling-based access protocols and orthogonal multiple access methods [1]. As networks evolve, the 6G vision outlines more stringent performance requirements for scenarios involving massive communication. This vision necessitates a connection density ranging from $10^{6}$ to $10^{8}$ devices per square kilometer[2]. As the number of users increases, dependence on orthogonal multiple access will lead to excessive overhead, substantially affecting communication efficiency.
To overcome the limitations of existing scheduling-based random access protocols, the grant-free non-orthogonal multiple access scheme has been proposed [1]. By eliminating the need for complex handshaking between users and the base station (BS), terminals can directly transmit uplink signals composed of non-orthogonal preamble sequences and payload data. This approach effectively reduces communication delays while enabling more devices to access the network simultaneously. However, it also increases the complexity of designing receiver detection algorithms. Previous studies have proposed several advanced receiver schemes, notably sourced random access and unsourced random access [3]. Sourced random access [4] employs pre-assigned preamble sequences to distinguish user identities and facilitates user activity detection and channel estimation using these preambles. In contrast, unsourced random access employs a common codebook, utilizing codewords to convey information and select coding, interleaving, and spreading techniques [5]. Moreover, reference [6] investigates a massive random access scheme utilizing individual user codebooks. This scheme recognizes the uniqueness of user codewords and encodes bits during the codeword selection process, integrating features from both sourced and unsourced random access mechanisms previously described. The reference [6] provides an in-depth performance analysis of massive random access under multiple-input multiple-output (MIMO) Rayleigh channels, taking into account the presence or absence of channel state information (CSI) at the receiver and whether the number of active users is known, thereby offering theoretical insights into performance limits. This paper mainly focuses on sourced random access schemes as in [4], tackling algorithm design challenges related to AUD, CE, and $D$. For simplicity, we continue to refer to the discussed approach as massive random access.
@@ I-A Related Work
Once the AUD and CE processes are completed, the massive MIMO $D$ problem at BS often becomes overdetermined. Various advanced detection algorithms have been developed to address this challenge. In addition to traditional methods like least squares (LS) and linear minimum mean square error (LMMSE), algorithms such as OAMP [18], expectation propagation (EP) [19], and deep learning-based techniques [20, 21, 22, 23] have demonstrated excellent performance. However, these $D$ methods rely on the CSI estimated in the earlier stage, and errors in CSI estimation will degrade the detection performance.
Performing joint CE and $D$ (JCEDD) can enhance overall achievable performance. The work [21] presents an iterative JCEDD scheme that accounts for channel estimation errors during data detection and employs a model-driven OAMP unfolding algorithm to improve detection. In the CE phase, demodulated data are used as pilots to augment LMMSE CE, resulting in superior performance. Additionally, several studies have addressed the joint AUD, CE, and $D$ problems, where message-passing algorithms or optimization-based methods have been developed to tackle these tasks [24, 25, 26]. Despite their strong performance, these algorithms require elaborate system configurations and dedicated algorithm design, which may not generalize effectively to practical systems or complex channel environments that are challenging to model, thus limiting their applicability. Consequently, they are not included in the comparisons of this paper.
@@ I-B Our Contributions
This paper proposes a new two-stage algorithm to solve the AUD, CE, and $D$ tasks in massive random access systems. The first stage identifies the active users set (AUS) through a Transformer-based detector, while the second stage utilizes generative diffusion models (GDMs) to perform JCEDD for active user equipment (UEs). The existing Transformer-based deep learning AUD scheme [10] fails to effectively address the challenges posed by variable pilot lengths in massive random access problems. Furthermore, the performance of current JCEDD schemes relying on ALD-based generative methods [33] has not been comprehensively evaluated and is susceptible to converging to local optima. To tackle these issues, our research makes the following key contributions:
@@
  • We model the CE and $D$ tasks as reverse diffusion processes that sample from a joint distribution. These processes gradually transform initial noise into the original channel and data symbols. By employing advanced predictor-corrector samplers, our scheme outperforms existing JCEDD schemes.
  • To accommodate the distinct characteristics of unknown continuous channel distributions and known discrete data constellation distributions, we propose an asynchronous iterative CE and $D$ framework. This framework integrates an LMMSE-based CE initialization method, which reduces the number of sampling steps and enhances convergence performance.
@@ III Problem Formulation
The proposed massive random access system is depicted in Fig. 1, where the BS needs to solve the problems of AUD and JCEDD. Specifically, we divide the AUD, CE, and $D$ tasks into the two stages. In the first stage, we use a single model to derive the AUS information from the received pilots of variable lengths. In the second stage, we perform the JCEDD using the powerful generation capability of diffusion models.
@@ III-B Stage 2-JCEDD
Different from typical passing algorithms that factorize the joint distribution (8) into a factor graph and then inferring the posterior distribution of each element through message passing, our proposed scheme directly samples the signals from the posterior distributions (9a) and (9b). This approach was first implemented in [27, 31, 32, 33], where the ALD sampler is used to sample from the posterior distribution in the CE or $D$ problem.
@@ IV-B Heterogeneous Transformer and Activity Decoder
Assuming that the heterogeneous Transformer network consists of NN encoder layers, we conduct activity estimation for $K$ users based on the K+1K\!+\!1 sequences output by the NN-th encoder layer. This task is performed by the activity decoder, which computes the heterogeneous MHA between the final output feature sequence of the received signal $\bm{z}_{K+1}^{N+1}$ and pilot sequences $\{\bm{z}_{k}^{N+1}\}_{k=1}^{K}$. The resulting MHA output, $\bm{z}_{c}$, represents the synthesized contributions of all pilot sequences to the received signal. Next, the user activity is estimated as the weighted correlation between $\bm{z}_{c}$ and each pilot feature sequence. It is given by:
@@ V Generative Diffusion Models Driven JCEDD Using Predictor-Corrector Sampler
Given the active UE set, we need to recover the channel and corresponding data symbols of the active UEs. Before introducing the proposed JCEDD schemes, we briefly review the relationship between existing generative models and their SDE descriptions. We then address the CE problem, detailing the calculation of posterior score functions within the predictor-corrector (PC) sampler framework. Subsequently, we extend the application of the PC sampler to the $D$ problem. Finally, we combine the CE and $D$ methods to propose an iterative algorithm for JCEDD and analyze its complexity.
@@ V-B3 Channel Sampling Process
Utilizing the derived prior score function and likelihood score function, we numerically solve the reverse SDE (V-B). Specifically, by dividing the interval $t\!\in\!(0,\,1]$ into NN discrete steps, indexed as $i\!=\!0,1,\dots,N-1$, we obtain the following iterative rule for the reverse diffusion sampler (which is also called predictor):
@@
where $\epsilon_{i}\!=\!2\sigma_{i}\big(r\|\bm{Z}\|_{2}/\|\nabla_{\tilde{\bm{H}}_{a}^{i}}\log p(\tilde{\bm{H}}_{a}^{i}|\bm{Y})\|_{2}\big)^{2}$ is the step size in the $i$-th iteration of the ALD algorithm [38]. Additionally, $r$ is an empirically chosen hyperparameter, set to $r=0.3$ in our simulations. Existing ALD-based studies on CE [27], $D$ [31], and JCEDD [33] can be regarded as corrector-only samplers with potential for performance improvement. Therefore, we address the JCEDD problem using the PC sampler illustrated in Fig. 4.
@@ V-C Data Detection
Upon reviewing the formulas (2) and (7), the $D$ problem can be characterized as $\bm{Y}_{d}\!=\!\bm{X}_{a}\bm{H}_{a}\!+\!\bm{W}_{d}$. By transposing the formula and representing the model in a real-valued form, the signal transmission model can be reformulated as:
@@
where $\lambda_{x}\!=\!2.5$ is an empirically chosen hyperparameter. The $D$ algorithm can also be derived based on PC-sampler. Due to space constraint, the detailed $D$ algorithm is omitted, but its basic iterative procedure is similar to that of Algorithm 1.
@@ V-D Overall Algorithm
In the CE and $D$ algorithms above, the data $\bm{X}_{a}$ is assumed to be known during CE, while the channel $\bm{H}_{a}$ is assumed to be known during $D$. In practical transmission, only the pilot $\bm{P}_{a}$ and the received signals $\bm{Y}$ are known. Therefore, we design an iterative algorithm to perform JCEDD. The proposed PC sampler-based JCEDD method is depicted in Fig. 5, and the algorithmic framework is summarized as Algorithm 2.
The issues we need to address are: 1) A typical SDE sampling algorithm starts by sampling from random noise, resulting in a high number of iterations. 2) In the existing ALD-based JCEDD [33], each CE iteration is followed by one $D$ iteration. Due to the different distributions of the channel and data symbols, their convergence speeds are inconsistent, causing the $D$ to fall into local optima before the completion of the CE iterations. We address these issues by the following improvements in the proposed sampling mechanism.
@@
    By reshaping the estimated $\hat{\bm{h}}_{a}$ into matrix $\hat{\bm{H}}_{a}$, the initial $D$ result can be obtained by applying zero forcing (ZF) to (27), i.e., $\hat{\bm{X}}_{a}\!=\!\bm{Y}_{d}\bm{H}_{a}^{\rm H}\big(\bm{H}_{a}\bm{H}_{a}^{\rm H}\big)^{-1}$.
  • We design an asynchronous alternating iteration mechanism. For channel datasets with complex distributions, the calculation of posterior score is more complicated than $D$ that relies solely on model-based iterations with closed-form expression. We update the $D$ results after every $N_{\text{update}}$ channel iterations (Line 6 in Algorithm 2). To ensure the optimal $D$ results under the current channel error, we conduct $N_{X}$ steps of PC sampler each time the data symbols are updated (Lines 9-12).
@@
1:Initialize: Acquire initial CE $\hat{\bm{H}}_{a}$ via LMMSE, initial $D$ $\hat{\bm{X}}_{a}$ via ZF, and start point $i^{*}$ via (33);
@@ VI-A2 Channel Prior Score Network Setup
During the testing phase, the reverse SDE process for CE is discretized into $N_{H}\!=\!1,500$ steps. Given that an LMMSE initialization is employed, only $i^{*}$ of the $N_{H}$ steps are executed. The reverse SDE process for $D$ is also discretized into $N_{X}\!=\!1,500$ steps, using a noise scheduling function defined as $\tau(t)\!=\!\tau_{\min}\big(\frac{\tau_{\max}}{\tau_{\min}}\big)^{t}$, where $\tau_{\text{max}}\!=\!1$ and $\tau_{\text{min}}\!=\!0.01$. Additionally, the hyperparameters are set to $\lambda_{h}\!=\!\lambda_{x}\!=\!2.5$.
@@ VI-A3 Evaluation Metrics
We utilize activity error probability (AEP), normalized mean square error (NMSE) and bit error rate (BER) to evaluate the performance of AUD, CE and $D$, respectively. These metrics are defined respectively as $\text{AEP}\!=\!\mathbb{E}\big\{\frac{1}{K}\sum_{k=1}^{K}|\alpha_{k}\!-\!\hat{\alpha}_{k}|\big\}$, $\text{NMSE}\!=\!\mathbb{E}\big\{\frac{\|\bm{H}_{a}\!-\!\hat{\bm{H}}_{a}\|^{2}}{\|\bm{H}_{a}\|^{2}}\big\}$ and $\text{BER}\!=\!\mathbb{E}\big\{\frac{E_{b}}{K_{a}L_{d}\log_{2}|\mathcal{X}|_{c}}\big\}$, where $E_{b}$ denotes the total number of error bits for active UEs.